OpenStudy (anonymous):

I need help with a simple radical expression real quick. fan and medal!!! just wait 30 seconds and ill have the expression in the comments

OpenStudy (anonymous):

whowww

OpenStudy (anonymous):

\[\sqrt{14q}\times 2 \sqrt{4q}\]

OpenStudy (anonymous):

woww what

OpenStudy (anonymous):

katra i already gave you medal for last question i had help with

OpenStudy (anonymous):



OpenStudy (anonymous):

is that 4 sqrt 14 sqrt?

OpenStudy (anonymous):

4 sqrt 14q^2?

OpenStudy (anonymous):

@katragaddasaichandra

OpenStudy (anonymous):

ya